At a major retail store, 25 percent of the employees use internet banking. For a randomly
selected group of five employees, calculate the probability that three use internet banking.
(6 Marks)
4.3 State four properties of the Normal Distribution function. (4 Marks)
4.4 An automatic machine fills jars of jam with a mean net weight of 340 grams. Assume a
normal distribution with a standard deviation of 8 grams. What is the probability that a
randomly selected jar of jam weighs between 338 grams and 344 grams.
1
Expert's answer
2019-04-30T09:11:20-0400

4.3.

  1. The normal curve is symmetrical about the mean μ;
  2. The mean is at the middle and divides the area into halves;
  3. The total area under the curve is equal to 1;
  4. It is completely determined by its mean and standard deviation.


4.4. P(338<X<344)=P((338-340)/8<Z<(344-340)/8)=P(-0.25<Z<0.5)=

=P(Z<0.5)-P(Z<-0.25)=P(Z<0.5)-1+P(Z<0.25)=

=0.6915-1+0.5987=0.2902.